OpenStudy (jjuden):

The height of a triangle is 4 m less than its base. The area of the triangle is 48 m^2. Find the length of the base.

OpenStudy (jjuden):

a 7 m b 8 m c 11 m d 12 m

OpenStudy (shamim):

let the base is=x

OpenStudy (shamim):

so the height is=x-4

OpenStudy (shamim):

the area of a triangle is 1/2(x)(x-4)=48

OpenStudy (shamim):

area of a triangle is =(1/2)*base*height

OpenStudy (anonymous):

There are a couple of ways to do this. Method 1: New width = 12+x New length = 14+x New area = 255 (12+x) (14+x) = 255 x^2 + 26x + 168 = 255 x^2 + 26x − 87 = 0 (x + 29) (x − 3) = 0, with x > 0 x = 3 New width = 12+3 = 15 m Method 2: Length is 2 more than width. If we add same amount to both, then length is still 2 more than width. w = width of larger garden w+2 = length of larger garden w (w+2) = 255 w^2 + 2w − 255 = 0 (w + 17) (w − 15) = 0, with w > 0 w = 15 New width = 15 m

OpenStudy (shamim):

@onlymath r u right with ur solution

OpenStudy (johnweldon1993):

I'm not sure what happened up there? but shamim's method was good... 1/2(x)(x - 4) = 48 Lets multiply everything by 2 to cancel out that 1/2 x(x - 4) = 96 Distribute the 'x' x^2 - 4x = 96 Now we can either plug that into the quadratic formula or just factor it...looks like factoring would be easy enough so x^2 - 4x - 96 = 0 a * b = -96 a + b = -4 Looks like the 2 numbers that multiply to make -96 AND add to make -4 are \(\large \space 12 \space \text{and} \space 8\) (x -12)(x + 8) So it looks like x will be either 12 or -8 (but we know it cannot be -8 ...so we arrive at 12 for your answer